OpenStudy (anonymous):

Find the surface area of the cone to the nearest square unit. Use pi = 3.14. The lateral area is 3cm and the radius is 6cm.

OpenStudy (anonymous):

Well if you know the lateral area then all you have to do it add it up with the area of the base.

OpenStudy (anonymous):

Since this is a cone the base will be a circle

OpenStudy (anonymous):

The area of a circles is: \[]\pi r^2\] where r= radius

OpenStudy (anonymous):

I am having problems getting this answer. The possible answers are

OpenStudy (anonymous):

please help me out with this!

OpenStudy (anonymous):

I gave you the formula for finding the area of a circle....where are you having problems?

OpenStudy (anonymous):

Btw you did not give me the lateral area

OpenStudy (anonymous):

First find the area of the circle

OpenStudy (anonymous):

\[area= \pi r^2\] \[\pi = 3.14 \]\[r=6cm\]

OpenStudy (anonymous):

ups didn't saw! the lateral of the area is 3cm

OpenStudy (anonymous):

lateral area*

OpenStudy (anonymous):

never mind I got it! thanks a lot for the help! it helped me a lot!